OpenStudy (anonymous):

Given the trinomial 2x2 + 4x - 2, what is the value of the discriminant?

OpenStudy (anonymous):

@SarahBrown @mathman806 @alejandra1995 @scott

OpenStudy (anonymous):

20

OpenStudy (anonymous):

The discriminant is the value of b^2 - 4ac. In this problem, a = 2, b = 4, and c = -2. So, go ahead, and compute the value of b^2 - 4ac.

OpenStudy (anonymous):

20 isnt in the answer box, I have 10, 0 32,36

OpenStudy (anonymous):

discriminant =b(square)-4ac =4-(4*2*-2) =4+16 =20

OpenStudy (anonymous):

hmmmmmm

OpenStudy (anonymous):

Croslali...we dont just give answers as the student doesnt learn anything. Show all work!

OpenStudy (anonymous):

Besides you gave the incorrect answer!

OpenStudy (anonymous):

Yeah and im in an online algebra2 class & im not quite getting this...

OpenStudy (anonymous):

sorry it is 32

OpenStudy (anonymous):

Thank you so much!!!!!

OpenStudy (anonymous):

Morgan...as I showed you before, the discriminant is the value of b^2 - 4ac...

OpenStudy (anonymous):

4ac means 4 times a times c

OpenStudy (anonymous):

discriminant =b(square)-4ac=16-(-16)=32

OpenStudy (anonymous):

Thank yall! Can I get yalls help pretty much whenever?

OpenStudy (anonymous):

Pretty much, yes!

OpenStudy (anonymous):

thank yall so much!